IF A BALLON WITH SAND IS COMING DOWN WITH UNIFORM VELOCITY.AFTER IT COMES DOWN HOW MUCH SAND SHOULD WE REMOVE FROM IT SO THAT IT GOES UP WITH UNIFORM VELOCITY?FORMULA?

Grade:

1 Answers

kshitij aggarwal
10 years ago

The various forces on the balloon are 

1) weight of sand + balloon ( always downwards)

2) air resistance  F  ( opposite to direction of motion)

3) Buoyant force  B  ( always upwards )

 

Say initial mass of balloon is m1.  So we can write

                       M1g = B + F

Now say the final mass of the balloon after removing some sand be M2.

     M2g + F = B

so M1 - M2 = 2F /g